OpenStudy (jonathanvaldes98):

are these both the same problem y = -3/2x + 4 –3x + 2y = 8

OpenStudy (akashdeepdeb):

No.

OpenStudy (jonathanvaldes98):

so whats diffrent. I have so solve it but i dont no much about the topic i odnt need the answer but if you could explain it.

OpenStudy (akashdeepdeb):

The first equation looks like it is the same but it isn't Look \(y = \frac{-3x}{2}+4\) Multiply 2 to both sides \(2y = -3x + 8\) Add 3x to both sides \(2y + 3x = 8\) THE SECOND EQUATION IS \(–3x + 2y = 8\) Understood how it is different? :D

OpenStudy (jonathanvaldes98):

took a while but i see it now lol thanks

OpenStudy (akashdeepdeb):

:)

OpenStudy (jonathanvaldes98):

one last question to get the y-intersept of the first one you would solve it right

OpenStudy (akashdeepdeb):

Yeah.

OpenStudy (jonathanvaldes98):

ok thanks

OpenStudy (akashdeepdeb):

.For y-intercept put the x value as 0. Because when x=0 then only the line intersects the y-axis.

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

No theyre not.... set youre x value at 0 to solve.